# Elements of World Knowledge (EWoK): A cognition-inspired framework for evaluating basic world knowledge in language models

### v1

@@ Abstract
The ability to build and leverage world models is essential for a general-purpose AI agent. Testing such capabilities is hard, in part because the building blocks of world models are ill-defined. We present Elements of World Knowledge (EWoK), a framework for evaluating world modeling in language models by testing their ability to use knowledge of a concept to match a target text with a plausible/implausible context. EWoK targets specific concepts from multiple knowledge domains known to be vital for world modeling in humans. Domains range from social interactions (help/hinder) to spatial relations (left/right). Both, contexts and targets are minimal pairs. Objects, agents, and locations in the items can be flexibly filled in enabling easy generation of multiple controlled datasets. We then introduce EWoK-core-1.0, a dataset of 4,374 items covering world knowledge domains. We evaluate open-weights large language models (B–B parameters) across a battery of evaluation paradigms along with a human norming study comprising 12,480 measurements. The overall performance of all tested models is worse than human performance, with results varying drastically across domains. These data highlight simple cases where even large models fail and present rich avenues for targeted research on LLM world modeling capabilities.
@@ 1 Introduction
Large language models (LLMs) acquire a substantial amount of knowledge from their training data, both through direct memorization and through learning co-occurrence-based text patterns (Bender and Koller, 2020; Grand et al., 2022; Pavlick, 2022). This knowledge comprises both knowledge about language (e.g. word meanings and rules of syntax) and knowledge about the world (e.g. social conventions and physical properties of objects). Contemporary LLMs’ knowledge of language is very robust (Mahowald et al., 2024), as evidenced by the fluency of the text that they generate. But how robust is their understanding of the basic social, physical, and relational concepts that are foundational to our everyday experience?
In this paper, we present Elements of World Knowledge (EWoK)111Data and associated code are available at: http://ewok-core.github.io, a flexible framework for evaluating world modeling in LLMs (see Figure 1 for an overview). The EWoK framework consists of: (a) several domains that constitute the foundation for basic human world knowledge and are processed by dedicated cognitive and neural systems; (b) specific concepts within each domain; (c) a set of item templates (and modular components to procedurally generate these templates) that test knowledge of these concepts by contrasting plausible and implausible context–target combinations; (d) a set of fillers to populate the templates, such that each template can be used multiple times; (e) a pipeline to generate a specific set of items (a dataset) based on these source materials.
@@
Our framework targets specific cognitive concepts (or concept pairs, such as friend/enemy). Concept knowledge is not limited to definitions, but can be used productively across a wide range of scenarios. Thus concepts leveraged in context are the first-class object of the EWoK framework, as opposed to individual sentences or facts. This approach stands in contrast to many NLP benchmarks, which often aim to evaluate knowledge based on individual items, or to develop sophisticated tasks that require drawing on many concepts simultaneously. Although real-world tasks can be more useful for comparing model performance, individual item complexity makes it hard to assess why a model fails. For instance, LLMs exhibit mixed performance on theory of mind tasks, which has led to disagreements about whether LLMs fail at inferring mental states or auxiliary cognitive abilities, such as knowing what it means for a box to be “transparent” (Kosinski, 2023; Ullman, 2023). Our framework mitigates this problem by explicitly linking the items with the concepts that they test.
@@
World knowledge is a notoriously fuzzy capability; which knowledge domains should one focus on? Here, we select a range of domains that have been shown to recruit dedicated cognitive and/or neural machinery in humans, such as knowledge of intuitive physics (McCloskey, 1983; Battaglia et al., 2013), knowledge of physical and spatial relations (Hafri and Firestone, 2021), intuitive number sense (Dehaene, 2011), social reasoning (Carlson et al., 2013), and reasoning about agents that involves both physical and social knowledge (Liu et al., 2024). These knowledge domains are not specific to language (Jackendoff, 2002); in fact, many are present in preverbal infants (Spelke and Kinzler, 2007). However, language contains a rich amount of information that reflects grounded world knowledge (Roads and Love, 2020; Abdou et al., 2021; Patel and Pavlick, 2021), such that LLMs might in principle acquire such domain-specific knowledge from text alone.
@@
To evaluate basic world modeling, we use combinations of plausible vs. implausible context–target pairs. Plausibility here serves as a proxy for factual accuracy: instead of deciding whether a factual statement is true or false, a model needs to determine whether a given scenario makes sense (is plausible). Having an accurate world model is necessary for consistently distinguishing plausible and implausible scenarios no matter how they are worded.
@@
Both contexts and targets in EWoK have a minimal-pairs design, such that a specific targeted change to a sentence (e.g., left right) results in an opposite result (plausible/implausible combination). This approach can help identify specific manipulations that LLMs are and are not sensitive to, with the goal of targeted diagnostics. It can also be leveraged in the future for mechanistic interpretability research.
@@
LLMs have a remarkable capacity for memorization, such that many plausible and implausible sentences can be distinguished solely based on their presence in the training data (e.g., The fox chased the rabbit is more common than The rabbit chased the fox). In contrast, our framework tests LLMs’ ability to evaluate contextual plausibility, such that the same exact target (The piano is left of Ali.) is either plausible or implausible depending on the context (see Figure 1 right).
Our paper is structured as follows. In Section 2, we review prior research related to this topic. In Section 3, we describe the core components of the EWoK framework. In Sections 4 and 5, we describe our evaluation strategy and the practical steps that we take to reduce the risk of dataset contamination. In Section 6, we show LLM performance on EWoK-core-1.0, a dataset generated via the EWoK framework. In Section 7, we discuss our results and future prospects for this line of work.
@@ 2 Related Work
The capabilities we evaluate are closely related to the notion of commonsense knowledge, an area that has numerous text-based benchmarks (e.g., Levesque et al., 2012; Sakaguchi et al., 2021; Zellers et al., 2019), including targeted evaluations of physical (Bisk et al., 2020) and social commonsense (Sap et al., 2019). LLMs often struggle on such commonsense benchmarks, likely due to the reporting bias in their training data (Shwartz and Choi, 2020): conversations and texts typically do not include commonly observed or obvious information (Gordon and Van Durme, 2013). Thus, although our items target basic world knowledge and are not designed to be challenging, LLMs might still struggle with them due to the reporting bias.
@@
Our evaluation is also related to works on natural language inference and entailment. The recognizing textual entailment (RTE) task Dagan et al. (2010) poses two sentences (a text expression and and hypothesis ) to a system and asks it to determine whether follows from . The natural language inference (NLI) task follows a similar challenge (Bowman et al., 2015; Williams et al., 2018; Conneau et al., 2018) and involves making a 3-way judgment about whether a premise entails, contradicts, or is neutral relative to a hypothesis. EWoK asks whether a target sentence is plausible given a context , which might—but does not have to—indicate an entailment relationship between the two. Though widely successful as a challenge, a large body of subsequent work has highlighted issues with RTE- and NLI-style evaluation: language models can often use heuristics (such as artifacts left behind by human annotators and lexical statistics) to “solve” the task without meaningful semantic understanding or reasoning (Poliak et al., 2018; Liu et al., 2020; McCoy et al., 2020; Gururangan et al., 2018). We address this limitation by (1) posing the task as a minimal pair where each of two targets must be preferred over its alternative given the right context, making reliance on target plausibility alone impossible; (2) annotating minimal pair contrast type to test whether which item design features drive model performance; and (3) testing the relationship between LLM performance and surface-level item properties, such as item length, average word frequency, and performance of a baseline bag-of-words embedding model.
@@
Our approach to dataset design is similar in spirit to the bAbi framework (Weston et al., 2016), which used simple synthetic tasks probing world knowledge and reasoning; however, our items are both simpler in design (they target individual concepts and do not, as of yet, require multi-chain reasoning or transitive inference) and harder in practice (a minimal pair, context-dependent design greatly reduces the availability of response heuristics, a serious problem in bAbi; Kaushik and Lipton, 2018).
@@
The minimal pair design is common in datasets inspired by psycholinguistics and cognitive science, such as SyntaxGym (Gauthier et al., 2020), BLiMP (Warstadt et al., 2020), and COMPS (Misra et al., 2023). In particular, it has previously been used to evaluate the models’ ability to distinguish plausible and implausible events (Pedinotti et al., 2021; Kauf et al., 2023), a task that draws heavily on commonsense knowledge. The popular Winograd Schema Challenge also had a minimal pairs setup (Levesque et al., 2012), even though that requirement was relaxed in later versions. We here extend this approach by employing a minimal pairs-of-pairs design, where both context and target sentences have a minimal pair counterpart.
@@
Until 2023, the dominant approach of assessing language model performance on a minimal pair has been to calculate each item’s (pseudo) log probability under the model. This method is effective at distinguishing grammatical and ungrammatical sentences (e.g., Warstadt et al., 2020), plausible and implausible events (Kauf et al., 2023), and relevant vs. irrelevant object properties (Misra et al., 2023), while often being calibrated to human sentence ratings and processing costs (Lipkin et al., 2023; Shain et al., 2024). Yet raw log probabilities reflect a number of surface-level properties of the input, such as word frequency (Kauf et al., 2023) and the number of possible paraphrases (Holtzman et al., 2021). An alternative approach, recently made possible with more powerful LLMs, is to prompt an LLM to to rate item plausibility, either absolute (on a Likert scale) or relative to the other item in the minimal pair. This approach can theoretically result in more task-specific estimates. However, for a range of linguistic and word prediction tasks, LLMs actually perform worse with direct prompting than via implicit log probability assessment (Hu and Levy, 2023), likely because of additional task demands imposed by the need to decipher instructions in the prompt (Hu and Frank, 2024). Thus, we report both log probability comparisons and explicit prompting results.
@@ 3 The Framework
We provide a flexible generative synthetic data pipeline (Figure 1), capable of producing many diverse datasets, each with unique specifications and statistics, while preserving metadata and decision traces. In Section 6, we use this framework to generate EWoK-core-1.0, a systematic, broad-coverage, context-sensitive world knowledge dataset containing 4,400 items.
Item format
@@
Effective use of world knowledge incorporates both access to a rich set of priors about the world’s structure and the ability to integrate this information on-the-fly with surrounding context. Thus, our items challenge models to leverage concepts in context. Each item consists of two minimal pair contexts (e.g., : The piano is in front of Ali. Ali turns left., : The piano is in front of Ali. Ali turns right) and two minimal pair target sentences (e.g., : The piano is right of Ali., : The piano is left of Ali.). The two target concepts are juxtaposed such that in any item, and . Thus, base target probabilities and cannot serve as plausibility cues: a model has to rely on context to establish plausibility.
Domains & concepts
EWoK is designed around domains of general world knowledge, with the current set shown in Figure 3. Each domain includes a set of concepts. For example, the domain social relations includes friend, enemy, teacher, student, boss and subordinate, and so on.
Dataset generation procedure
Each concept is associated with several items that test knowledge of the concept (often, but not always, by contrasting it with another concept). Items are created in a flexible yet controlled manner using the EWoK dataset generation procedure. At the core of the framework, we have atomic units and combination rules that, subject to constraints, lead to the generation of templates which are then populated with fillers—this leads to generating many more carefully-controlled items than a single-step template-filling approach could generate.
Contexts & targets
@@
A target is a simple sentence that incorporates a concept; each domain has one or more associated targets depending on the lexical properties of the concepts in that domain. A contrasting target pair is generated using one of two mechanisms: concept swap, which contrasts the same target with different concepts filled in, and variable swap, which swaps two objects or agents mentioned in the target (only possible for certain concepts). For instance, {agent1} is to the left of {agent2} can be contrasted with {agent1} is to the right of {agent2} (concept swap) or with {agent2} is to the left of {agent1} (variable swap).
A context pair is one or more minimal pair(s) of sentences that can each be paired with a target pair such that but not matches and but not matches . They are typically specific to an opposing concept pair (left/right) or to a single concept (left; only possible when variable swap is allowed). Similarly, a contrasting context pair is generated using one of two mechanisms, namely, a filler swap and a variable swap. A filler swap uses contrasting fillers to generate two versions of the item. A variable swap changes positions of two entities of the same kind to change the nature of the relationship or dynamic between them as long as the concept supports it.
Templates & fillers
Each collection of concepts, contexts, and targets can then be compiled to a set of templates, partial items with typed variables describing the range of fillers for which the intended meaning holds. For example, in the template {object2:can_bounce=True} bounced off {object1} from below, any object may fill {object1}, e.g., the desk or the crate, whereas {object2:can_bounce=True} must be filled by an object marked with a flag for can_bounce=True, e.g., the ball or the tire. Across domains we list over filler items across classes with type restrictions.
@@
When populating the templates with fillers, users can specify various custom parameters, e.g.: (1) specify the number of items to generate from each template; each full set of items is referred to as a version; (2) select whether fillers should be held constant across all items in a version or allowed to vary; and (3) apply transformations to filler restrictions at compile-time, for instance restricting all agents to take non-Western names via agent->agent:western=False, or swapping all objects with nonce words via object->nonword. Such flexibility allows for controlled experimentation of the features modulating model performance. With these arguments and others easily accessible from the command line interface, users are supported to generate an enormous space of EWoK variations, and may easily add new domains, concepts, fillers, and more.

### v2

@@ Abstract
The ability to build and reason about models of the world is essential for situated language understanding. But evaluating world modeling capabilities in modern AI systems—especially those based on language models—has proven challenging, in large part because of the difficulty of disentangling conceptual knowledge about the world from knowledge of surface co-occurrence statistics. This paper presents Elements of World Knowledge (EWoK), a framework for evaluating language models’ understanding of the conceptual knowledge underlying world modeling. EWoK targets specific concepts from multiple knowledge domains known to be important for world modeling in humans, from social interactions (help, deceive) to spatial relations (left, right). Objects, agents, and locations in the items can be flexibly filled in, enabling easy generation of multiple controlled datasets. We then introduce EWoK-core-1.0, a dataset of 4,374 items covering world knowledge domains. We evaluate open-weights large language models (B–B parameters) and compare them with human performance. All tested models perform worse than humans, with results varying drastically across domains. Performance on social interactions and social properties was highest and performance on physical relations and spatial relations was lowest. Overall, this dataset highlights simple cases where even large models struggle and presents rich avenues for targeted research on LLM world modeling capabilities.
@@ 1 Introduction
Large language models (LLMs) acquire a substantial amount of knowledge from their training data (Bender and Koller, 2020; Grand et al., 2022; Pavlick, 2022). This knowledge comprises both knowledge about language (e.g. word meanings and rules of syntax) and knowledge about the world (e.g. social conventions and physical properties of objects). Contemporary LLMs demonstrate substantial knowledge of language (Mahowald et al., 2024), as evidenced by the fluency of the text that they generate. But how robust is their understanding of the basic social, physical, and relational concepts that are foundational to our everyday experience?
In this paper, we present Elements of World Knowledge (EWoK)111Data and associated code can be accessed at http://ewok-core.github.io , a flexible framework for evaluating world modeling in LLMs (Figure 1). The EWoK framework consists of: (a) several domains that constitute the foundation for basic human world knowledge and are processed by dedicated cognitive and neural systems; (b) a set of concepts within each domain; (c) a set of templates (and modular components to procedurally generate these templates) that test knowledge of specific concepts by contrasting plausible and implausible context–target combinations; (d) a set of fillers to populate the templates, such that each template can be used multiple times; (e) a pipeline to generate a specific set of items (a dataset) based on these source materials; and (f) an evaluation procedure that measures language models’ ability to distinguish plausible and implausible items.
@@
Our framework targets specific cognitive concepts (or concept pairs, such as left/right). Concept knowledge is not limited to memorized definitions, but can be used productively across a wide range of scenarios. Thus concepts (not isolated sentences or facts) are the primary focus of the EWoK framework. This approach stands in contrast to many NLP datasets, which often include naturalistic, lengthy stimuli that draw on multiple unspecified concepts. Although naturalistic tasks are useful for evaluating model performance in real-life settings, individual item complexity and lack of controls can make it difficult to assess why a model fails. For instance, LLMs exhibit mixed performance on theory of mind tasks, which has led to disagreements about whether LLMs fail at inferring mental states or auxiliary cognitive abilities, such as knowing what it means for a box to be “transparent” (Kosinski, 2024; Ullman, 2023). Our framework mitigates this problem by explicitly linking individual items with the specific concepts that they test.
@@
World knowledge is a notoriously fuzzy capability; which knowledge domains should one focus on? Dataset design in NLP is often centered around data availability, with examples sourced from video captions, Wikipedia articles, or social media logs. In contrast, here, we systematically select a range of knowledge domains that have been shown to recruit dedicated cognitive and/or neural machinery in humans, such as knowledge of intuitive physics (McCloskey, 1983; Battaglia et al., 2013), knowledge of physical and spatial relations (Hafri and Firestone, 2021), intuitive number sense (Dehaene, 2011), social reasoning (Carlson et al., 2013; Thomas, 2024), and reasoning about agents that involves both physical and social knowledge (Liu et al., 2024). These knowledge domains are not specific to language (Jackendoff, 2002); in fact, many are present in preverbal infants (Spelke and Kinzler, 2007). At the same time, traces of these signals are deeply embedded in our daily communicated language. Text itself thus contains rich information that reflects grounded world knowledge (Roads and Love, 2020; Abdou et al., 2021; Patel and Pavlick, 2021), and it may be reasonable to predict that LLMs would acquire this domain-specific knowledge from text alone.
@@
Concept understanding is often fuzzy. Thus, deciding whether a factual statement is true or false in isolation is often ill-defined. Instead, we use combinations of plausible vs. implausible context–target pairs. In EWoK, a model needs to determine which scenarios make more sense (more plausible). Having an accurate world model would enable a model to consistently distinguish plausible and implausible scenarios no matter how they are worded.
@@
Both contexts and targets in EWoK have a minimal-pairs design, such that a specific targeted change to a sentence (e.g., left right) yields an opposite effect (e.g., plausible implausible). This approach can help identify specific manipulations that LLMs are and are not sensitive to, with the goal of targeted diagnostics. Such controlled manipulations are also particularly well-suited for mechanistic interpretability research, which may seek to explain the circuits through which models deploy world knowledge in context.
@@
LLMs have a remarkable capacity for memorization, such that many plausible and implausible sentences can be distinguished solely based on their presence in the training data (e.g., The fox chased the rabbit is more common than The rabbit chased the fox). In contrast, our framework tests LLMs’ ability to evaluate contextual plausibility, such that the same exact target (The piano is left of Ali) is either plausible or implausible depending on the context (see Figure 1 right). Establishing this control is critical for isolating contextually-sensitive world knowledge from heuristic knowledge associated with surface forms.
Our paper is structured as follows. In Section 2, we review prior research related to the topics explored in this work. In Section 3, we describe the core components of the EWoK framework. In Section 4, we describe our evaluation strategy. In Section 5, we show LLM performance on EWoK-core-1.0, a dataset generated via the EWoK framework. In Section 6, we discuss our results and future prospects for this line of work. Finally, in Section 7, we discuss dataset release considerations.
| Domain | Templates | Concepts | Example | ||
| Social Interactions | 165 | 16 |
| Social Properties | 185 | 16 |
| Social Relations | 785 | 15 |
| Physical Interactions | 280 | 20 |
| Physical Dynamics | 75 | 12 |
| Physical Relations | 435 | 20 |
| Material Dynamics | 780 | 21 |
| Material Properties | 125 | 18 |
| Agent Properties | 1130 | 22 |
| Quantitative Properties | 195 | 18 |
| Spatial Relations | 245 | 14 |
@@ 2 Related Work
The capabilities we evaluate are closely related to the notion of commonsense knowledge, an area that has numerous text-based benchmarks (e.g., Levesque et al., 2012; Sakaguchi et al., 2021; Zellers et al., 2019), including targeted evaluations of physical (Bisk et al., 2020) and social commonsense (Sap et al., 2019). LLMs often struggle on such commonsense benchmarks, likely due to the reporting bias in their training data (Shwartz and Choi, 2020): conversations and texts typically do not include commonly observed or obvious information (Gordon and Van Durme, 2013). Thus, although our items target basic world knowledge and are not designed to be challenging, it is possible that LLMs could still struggle with them due to the reporting bias.
@@
Our evaluation is also related to works on natural language inference and entailment. The recognizing textual entailment (RTE) task Dagan et al. (2010) poses two sentences to a system (a text expression and hypothesis ) and asks that it determine whether follows from . The natural language inference (NLI) task follows a similar challenge (Bowman et al., 2015; Williams et al., 2018; Conneau et al., 2018) and involves making a 3-way judgment about whether a premise entails, contradicts, or is neutral relative to a hypothesis. EWoK asks whether a target sentence is plausible given a context , which might—but does not have to—indicate an entailment relationship between the two. Though widely successful as a challenge, a large body of subsequent work has highlighted issues with RTE- and NLI-style evaluation: language models can often use heuristics (such as artifacts left behind by human annotators and lexical statistics) to “solve” the task without meaningful semantic understanding or reasoning (Poliak et al., 2018; Liu et al., 2020; McCoy et al., 2020; Gururangan et al., 2018). We address this limitation by (1) posing the task as a minimal pair, where each of two targets is held static while paired contexts are used to modulate preferences, rendering it impossible to rely solely on target plausibility; (2) annotating minimal pair contrast type to test whether any item design feature drives model performance; and (3) testing the relationship between LLM performance and surface-level item properties, such as item length, average word frequency, and performance of a baseline bag-of-words embedding model.
@@
Our approach to dataset design is similar in spirit to the bAbi framework (Weston et al., 2016), which used simple synthetic tasks probing world knowledge and reasoning; however, our items are both simpler in design (they target individual concepts and do not require multi-chain reasoning) and are harder in practice (a minimal pair, context-dependent design greatly reduces the availability of response heuristics, a serious problem in bAbi; Kaushik and Lipton, 2018).
@@
The minimal pair design is common in datasets inspired by psycholinguistics and cognitive science, such as SyntaxGym (Gauthier et al., 2020), BLiMP (Warstadt et al., 2020), and COMPS (Misra et al., 2023). In particular, it has previously been used to test models’ ability to distinguish plausible and implausible events (Pedinotti et al., 2021; Kauf et al., 2023), a task that draws heavily on commonsense knowledge. The popular Winograd Schema Challenge also had a minimal pairs setup (Levesque et al., 2012), even though that requirement was relaxed in later versions. We here extend this approach by employing a minimal pairs-of-pairs design, where both context and target sentences have a minimal pair counterpart.
@@
Before prompt-based evaluation became commonplace, the dominant approach for assessing language model performance on a minimal pair has been to calculate each item’s (pseudo) log probability under the model. This method is effective at distinguishing grammatical and ungrammatical sentences (e.g., Warstadt et al., 2020), plausible and implausible events (Kauf et al., 2023), and relevant vs. irrelevant object properties (Misra et al., 2023), while often being calibrated to human sentence ratings and processing costs (Lipkin et al., 2023; Shain et al., 2024). Yet raw log probabilities reflect a number of surface-level properties of the input, such as word frequency (Kauf et al., 2023) and the number of possible paraphrases (Holtzman et al., 2021). An alternative approach, recently made possible with more powerful LLMs, is to prompt an LLM to rate item plausibility, either absolute (on a Likert scale) or relative to the other item in the minimal pair. This approach can theoretically result in more task-specific estimates. However, for a range of linguistic and word prediction tasks, LLMs actually perform worse with direct prompting than via implicit log probability assessment (Hu and Levy, 2023), likely because of additional task demands imposed by the need to decipher instructions in the prompt (Hu and Frank, 2024). Thus, we report both log probability comparisons and explicit prompting results.
@@ 3 The Framework
We provide a flexible generative synthetic data pipeline, capable of producing many diverse datasets, each with unique specifications and statistics, while preserving metadata and decision traces. Here, we discuss the design principles that make it possible to reuse the EWoK framework for systematically generating various world knowledge datasets. In Section 5, we use this framework to generate EWoK-core-1.0, a systematic, broad-coverage, context-sensitive world knowledge dataset containing items that probe concepts (for a full concept list, see Table A3).
Items test concept knowledge in context.
@@
Effective use of world knowledge incorporates both access to a rich set of priors about the world’s structure and the ability to integrate this knowledge on-the-fly with information about the current environment. Thus, our items challenge models to leverage concepts in context. Each item consists of two minimal pair contexts (e.g., : The piano is in front of Ali. Ali turns left., : The piano is in front of Ali. Ali turns right) and two minimal pair target sentences (e.g., : The piano is right of Ali., : The piano is left of Ali.). The two target concepts are juxtaposed such that in any item, and . Thus, base target probabilities and cannot serve as plausibility cues: a model has to rely on context to establish plausibility.
Concepts comprise domains of world knowledge.
EWoK is designed around domains of general world knowledge, with the current set of 11 domains shown in Table 1. We selected a range of domains that have been shown to recruit dedicated cognitive and/or neural machinery in humans, backed by extensive literature in cognitive science (introduced in Sec. 1, “Why cognition inspired?”). Domains were contributed based on past literature during the development of EWoK by a team of experts in the field (authors of this paper: professors, post-docs, and grad students in cognitive science and neuroscience across a few institutions). Each domain includes a set of concepts, also curated by the team. For example, the domain social relations includes friend, enemy, teacher, student, boss, subordinate, and others. The number of concepts in each domain is mentioned in Table 1.
Items are generated from concept-specific templates.
Each concept is associated with several items that test knowledge of the concept (often, but not always, by contrasting it with another concept). Items are generated from templates, which contain placeholders for specific objects, agents, and locations. The placeholder value should not affect the resulting plausibility judgments; thus, a template can be populated with an arbitrarily large number of fillers, which enables generating many carefully controlled items.
Templates consist of contrasting context and target pairs.
A template includes 2 targets and 2 contexts. A target is a simple sentence that incorporates a concept. A context is a short sequence of words that can be paired with a target to yield either a plausible or an implausible combination. Context and target pairs are designed in such a way that but not matches and but not matches . The differences between the two targets or the two contexts result in specific types of contrasts, shown in Figure 2.
@@
A contrasting target pair is generated using one of two mechanisms: concept swap, which contrasts the same target with different concepts filled in, and variable swap, which swaps two objects or agents (only possible for certain targets). For instance, AGENT-1 is AGENT-2’s teacher can be contrasted with AGENT-1 is AGENT-2’s student (concept swap) or with AGENT-2 is AGENT-1’s teacher (variable swap).
A contrasting context pair is generated using antonyms, negation, or variable swap.
Finally, the templates themselves test concept knowledge in either direct or indirect way. Direct tests rely on immediate interpretation of a concept, whereas indirect tests leverage the probabilistic relationship between concept and target ( doesn’t have to be true given but it’s more likely than given ).
Templates enable flexible yet controlled generation of items using typed fillers.
Placeholders in templates (e.g., OBJECT or AGENT) can be restricted to only allow fillers of specific types. For example, in the template {object2:can_bounce=True} bounced off {object1} from below, {object2:can_bounce=True} must be filled by an object marked with a flag for can_bounce=True, e.g., the ball. These type restrictions prevent generation of semantically anomalous or incomprehensible sentences (e.g., The blanket bounced off the table from below). Type restrictions were specified by the author team during the template development stage and later validated via human behavioral ratings (Section 4.4 and Appendix A.2).
To create an initial set of object and location fillers, we generated a set of arbitrary fillers based on constraints specified in templates (e.g., objects that can bounce, that have a smell) that was filtered down to ensure the ‘naturalness’ of the items that include these fillers. The agent fillers were selected from the list of culturally diverse high frequency male, female and non-gendered baby names. Across domains we list over filler items across classes with type restrictions. The full set of fillers is provided in the paper repo.
@@
When populating the templates with fillers, users wanting to use the EWoK framework to construct their own instance of an EWoK-core-1.0-like dataset can specify several parameters, e.g.: (1) the number of items to generate from each template; each full set of items is referred to as a version (corresponding to a random seed); (2) whether fillers should be held constant across all items in a version or allowed to vary; (3) whether to apply transformations to filler restrictions at compile-time, for instance restricting all agents to take non-Western names via agent->agent:western=False, or swapping all objects with nonce-words via object->nonword. Such flexibility allows for controlled experimentation of the features modulating model performance. With these arguments and others easily accessible from the command line interface, users are supported to generate many EWoK variations, and can additionally extend EWoK to add new domains, concepts, and fillers.